# XAMI - A Benchmark Dataset for Artefact Detection in XMM-Newton Optical Images

### v2

@@ 3 Results
Following recommendations in [kirillov2023segment], we utilise the focal loss and dice loss in a 20:1 weighted scheme. At this stage, predicted and actual masks can be directly compared. Unlike usual SAM implementations, we choose to train the Intersection-over-Union (IoU) head to provide more representative mean Average Precision (mAP) metrics (see Eq. 1). Also, when generating masks, we configure the model to allow three predicted mask outputs and select the final mask based on the highest IoU score. The overall loss calculation integrates both segmentation and IoU loss.
After training the YOLO and SAM models separately to optimise their individual performances, we freeze the YOLO layers, couple its predicted bounding boxes to the SAM Prompt Encoder and continue training the SAM Mask Decoder to refine the segmentation process for 10 additional epochs. The alignment of predicted and ground truth masks is managed using the Kuhn-Munkres assignment algorithm [https://doi.org/10.1002/nav.3800020109] by minimizing the IoU cost matrix. Due to higher spatial complexity of certain classes, particularly SL and Other, we select YOLO masks for faint objects of such classes at background level, as these predictions are more stable for low-intensity artefacts. We provide the segmentation mAP (see Table 3) using a fixed seed for reproducibility. The mAP formula for instance segmentation is given by:
where is the number of classes, and is the average precision for the -th class, calculated as the area under the precision-recall curve at different IoU thresholds.
| Category | mAP50 | mAP75 | mAP50-90 |
| Overall | 79.1 | 63.6 | 46.3 |
| Small | 68.7 | 50.6 | 39.3 |
| Medium | 80.0 | 63.7 | 45.3 |
| Large | 80.8 | 66.0 | 53.5 |
| CR | 96.8 | 93.4 | 64.7 |
| SR | 87.7 | 70.8 | 46.2 |
| SL | 84.6 | 66.7 | 50.9 |
| ROS | 60.4 | 34.7 | 33.3 |
| Other | 66.0 | 52.6 | 36.1 |

### v3

@@ 3 Results
Following recommendations in [kirillov2023segment], we utilise the focal loss and dice loss in a 20:1 weighted scheme. At this stage, predicted and actual masks can be directly compared. Unlike usual SAM implementations, we choose to train the Intersection-over-Union (IoU) head to provide more representative segmentation metrics. Also, when generating masks, we configure the model to allow three predicted mask outputs and select the final mask based on the highest IoU score. The total loss integrates both segmentation and IoU loss.
After training the YOLO and SAM models separately to optimise their individual performances, we freeze the YOLO layers, couple its predicted bounding boxes to the SAM Prompt Encoder and continue training the SAM Mask Decoder to refine the segmentation process for 10 additional epochs. The alignment of predicted and ground truth masks is managed using the Kuhn-Munkres assignment algorithm [https://doi.org/10.1002/nav.3800020109] by minimizing the IoU cost matrix. Due to higher spatial complexity of certain classes, particularly SL and Other, we select YOLO masks for faint objects of such classes at background level, as these predictions are more stable for low-intensity artefacts. We provide the precision and recall metrics (see Table 3) using a fixed seed for reproducibility. The formulas for precision and recall are defined as follows:
where TP - True Positive, FP - False Positive and FN - False Negative predictions.
Category | Precision | Recall |
Overall | 84.3 | 72.1 |
CR | 89.3 | 94.0 |
SR | 80.6 | 85.6 |
SL | 80.5 | 74.1 |
ROS | 71.1 | 73.3 |
Other | 100.0 | 33.3 |